Skip to main contentSkip to solution

If XX is a random variable which can assume values 0,1,2,30, 1, 2, 3 or 44 such that P(X=1)=P(X=2)P(X = 1) = P(X = 2) and 3P(X=3)=4P(X=4)=P(X=0)=183P(X = 3) = 4P(X = 4) = P(X = 0) = \frac{1}{8}, then P(X>0)P(X > 0) is:

Solution

✅ Correct Option: 4

To find P(X>0)P(X > 0), we can use the fundamental property that the sum of all probabilities in a probability distribution must equal 11.

  • P(X=0)=18P(X = 0) = \frac{1}{8}
  • 4P(X=4)=18  ⟹  P(X=4)=1324P(X = 4) = \frac{1}{8} \implies P(X = 4) = \frac{1}{32}
  • 3P(X=3)=18  ⟹  P(X=3)=1243P(X = 3) = \frac{1}{8} \implies P(X = 3) = \frac{1}{24}

The event X>0X > 0 includes the outcomes X=1,X=2,X=3,X=1, X=2, X=3, and X=4X=4. The complement of this event is X=0X = 0.

The formula for the complement is:

P(X>0)=1−P(X=0)P(X > 0) = 1 - P(X = 0)

Substituting the known value for P(X=0)P(X = 0):

P(X>0)=1−18P(X > 0) = 1 - \frac{1}{8}

P(X>0)=88−18P(X > 0) = \frac{8}{8} - \frac{1}{8}

P(X>0)=78P(X > 0) = \frac{7}{8}

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question